# Submitted to Phys. Rev. B. Excitonic Insulator and the Extended Falicov–Kimball Model Away from Half-Filling

### v1

@@ Abstract
We consider an extended spinless Falicov–Kimball model at an arbitrary doping level, focusing on the range of parameter values where a uniform excitonic insulator is stabilised at half-filling. We compare the properties of possible uniform phases and construct the Hartree–Fock phase diagrams, which include sizeable phase separation regions. It is seen that the excitonic insulator can appear as a component phase in a mixed-phase state in a broad interval of doping levels. In addition, in a certain range of parameter values the excitonic metal (doped excitonic insulator) is identified as the lowest-energy uniform phase. We suggest that this phase, which is unstable with respect to phase separation, may be stabilised when the phase separation is suppressed by the long-range Coulomb interaction. Overall, we find that excitonic correlations can affect the behaviour of the system relatively far away from half-filling.
@@ II THE MODEL AND MEAN-FIELD EQUATIONS
The self-consistent mean-field equations for the excitonic phase are derived in a standard way and take form
@@
$n_{c}-n_{d}$ $=$ $\frac{1}{N}\sum_{\vec{k}}\frac{\left(\xi_{{\vec{k}}}+t^{\prime}\epsilon_{\vec{k}}\right)\left(n^{1}_{\vec{k}}-n^{2}_{\vec{k}}\right)}{2\sqrt{(\xi_{{\vec{k}}}+t^{\prime}\epsilon_{\vec{k}})^{2}+4|U\Delta|^{2}}}\,,$ (4)
which is valid at all values of bandfilling $20<n<$ 2,
@@
$\epsilon^{1,2}_{\vec{k}}$ $=$ $\frac{1}{2}\left[E_{d}+Un+\left(1+t^{\prime}\epsilon_{k}\right)\right]\mp$ (6)
@@
are quasiparticle energies in the two new bands, $\epsilon_{\vec{k}}=-\cos k_{x}-\cos k_{y}(-\cos k_{z})$ is the tight-binding dispersion in two (three) dimensions, and $\xi_{\vec{k}}=E_{d}+U(n_{c}-n_{d})-\epsilon_{\vec{k}}$. The net energy of the excitonic phase can be evaluated as [cf. Eq. (2)]
@@
At half-filling and at $T\rightarrow 0$, the quantity $n^{2}_{\vec{k}}$ vanishes, while $n^{1}_{\vec{k}}\equiv 1$ for all values of $\vec{k}$; Eq. (7) then yields $E_{EI}$, the energy of EI. We shall assume [without loss of generality – see Eq. (3)] that the quantity $\Delta$ is real and positive.
@@
At smaller $U$, one also finds semimetal solutions with two partially-filled Hartree bands, centred around $Un_{d}$ (the broad band) and $E_{d}+Un_{c}$. Writing $\mu=Un_{d}+\lambda_{c}$, we then find for the energy difference $|t^{\prime}|\lambda_{d}$ between $\mu$ and the centre of the narrow band,
@@
[…]
@@
[…]
The rest of the paper is concerned with numerical analysis of these solutions and of the ensuing phase diagrams in the interaction range $|t^{\prime}|\ll U/2d\stackrel{{\scriptstyle<}}{{\sim}}1$ (weak to moderately strong coupling). We shall be interested in the low-temperature limit, $T\rightarrow 0$ throughout. While we consider the two-dimensional case only, it is expected that the mean-field analysis of a three-dimensional system would yield qualitatively similar results.
@@ III THE UNIFORM SOLUTIONS AND THEIR INSTABILITIES
We begin with the excitonic phase, which at half-filling, $n=1$, corresponds to the well-known excitonic insulator (EI) state, characterised by a non-zero value of the off-diagonal average $\Delta$ and a gap of the order of $U\Delta$ in the electron spectrum. Properties of EI phase and its stability have been extensively addressed in the literature[2, 3, 7]. Presently we find, that once the EI solution is present at $n=1$, a similar solution to the mean field equations, Eqs. (34) persists also in a certain range of doping values around $n=1$. The chemical potential then lies within the upper or lower hybridised band, away from the excitonic band gap, hence such a phase should be more properly called excitonic metal (EM). This term, which was used previously in the context of doped Mott insulators[15], is now more commonly applied to the case of doped and/or otherwise imperfect EI, with a non-zero density of states at the Fermi level[16, 18, 17].
Fig. 1 illustrates the typical behaviour of EM solution in the case of moderate (a) or weak (b) Coulomb repulsion $U$. Numerical calculations are somewhat tedious, in particular because even relatively small values of $t^{\prime}$ strongly affect the EM bandstructure and may, e.g., shift the minimum of the upper hybridised band away from $\vec{k}=0$.
@@
For $U=2$, the excitonic metal solution is found within the doping range $1.380.93<n<$ 1.38 and with increasing $n$ the value of $n_{d}$ increases from zero to one. Accordingly, at the two endpoints the excitonic solution merges with the two different single-band phases. At these endpoints, the off-diagonal average $\Delta$ shows square-root features; elsewhere, it has a maximum at $n=1$ and a smoother feature near $n=1.3$, the latter reflecting a feature in the bandstructure. In the weak-interaction regime of $U=0.5$, the excitonic metal solution arises for $1.120.91<n<$ 1.12, merging with the semimetal solution at the endpoints.
Importantly, everywhere away from half-filling the compressibility $\partial\mu/\partial n$ is negative, signalling an instability with respect to an inhomogeneity formation and ultimately to phase separation. The dependence $\mu(n)$ shows a jump at $n=1$. This point, where the value of $\mu$ lies within the gap between the lower (filled) and the upper (empty) bands [see Eq. (6)], has to be considered separately.
@@
We remark that analysis of a “pure” Falicov – Kimball model [Eq. (1) with $t^{\prime}=0$] yields qualitatively similar behaviour of mean-field solutions, although quantitative changes associated with the non-zero $t^{\prime}$ are in some cases appreciable. We include a finite $t^{\prime}<0$ in our analysis due to the peculiar features of the $t^{\prime}=0$ case. These include an instability of excitonic insulator at $n=1$ (for all values of $U$ and $E_{d}$), as clarified in Refs. [6, 7]. For the numerical data shown in the plots, we chose the values of $-t^{\prime}$ well above the respective critical values required to stabilise the homogeneous excitonic insulator state at half-filling; these critical values correspond to a second-order transition into a spatially modulated phase[6, 7].
@@
where the values of $V_{0}$ and/or $V_{1}$ exceed certain numerically small critical values. While these leads to numerical changes, we tentatively verified that at the qualitative level, the conclusions of this section (that the excitonic metal state has a negative compressibility and also, typically, the lowest energy) remain valid. The important difference is that in the presence of a bare hybridisation one has to distinguish not between non-hybridised ($\Delta=0$) and hybridised (excitonic metal or insulator) solutions as above, but rather between phases with small and large $\Delta$. In the former case, $\Delta$ vanishes in the limit $V_{0,1}\rightarrow 0$, as the corresponding solution evolves into either single-band or semimetal one. This does not happen in the case of large-$\Delta$ solutions, corresponding to EM or EI (notwithstanding an instability at small $V_{0,1}$, which is similar to the instability at small $t^{\prime}$). Throughout the rest of the paper, we will be including the effects of $t^{\prime}$ only.
@@ IV PHASE SEPARATION
At smaller $U$, this can preempt the EI – single-band PS1 phase separation and may also render the EI state at $x=1$ unstable (basically, the single-band state that would have been in thermodynamic equilibrium with the EI becomes unstable with respect to this second type of phase separation). We will now follow these two phase separation scenarios in a typical situation (the details may vary, depending on parameter values).
@@
With lowering $n$ further, the stability of the uniform phase $A$ is lost at a point $n=n_{A}^{*}$ where either $F_{A}$ or $F_{AB}$ vanishes (whichever occurs first). In the case where $F_{A}(n_{A}^{*})=0$ [while $F_{AB}(n_{A}^{*})<0$], lowering $n$ further leads to phase separation into phase A (with $n_{A}=n_{A}^{*}$) and the EI phase; the EI fraction expands until phase A disappears at half-filling ($n=1$) and the system turns into a uniform EI. At this point, the value of the chemical potential suffers a negative jump (within the EI spectral gap), and reducing $n$ further leads to a phase separation into EI and a single-band phase with empty narrow band.
@@
These two situations are exemplified by Fig. 6. For the values of parameters used in Fig. 6 (a), Eq. (15) is satisfied at $n<1.68$, and $F_{A}(n)$ changes sign at $n_{A}^{*}\approx 1.522$, signalling phase separation into a half-filled EI and the single-band phase A with $n=n_{A}^{*}>1$ (phase separation of the type PS1). On the other hand, the value of $F_{AB}$, which would describe phase separation into two different single-band phases, changes sign at $n\approx 1.492<n_{A}*$, and this latter scenario is therefore irrelevant (as the uniform phase A at this doping level is already unstable with respect to the other type of phase separation). In the case of Fig. 6 (b), Eq. (13) for those parameters values can be solved only in a very narrow doping range $1.6121.608<n<$ 1.612 [cf. Fig. 3 (b), where the overall negative slope of the dashed line in the centre of the figure is very small]. The value of $F_{AB}(n)$ changes sign at $n_{A}^{*}\approx 1.611$, whereas $F_{A}(n)$ vanishes at $n\approx 1.610<n_{A}^{*}$ and the EI state is therefore irrelevant. Phase separation is thus of the PS2 type, corresponding to a mixture of two single band phases with $n=n_{A}^{*}$ and $n=n_{B}^{*}\approx 0.756$. We note that the precise boundary between the two scenarios depends also on the value of $t^{\prime}$, as increasing the latter tends to tilt the balance in favour of PS1.
@@ V THE PHASE DIAGRAMS
Figure 7: (colour online) Calculated phase diagrams for the EFKM with $U=4$ and $t^{\prime}=-0.3$ (a), $U=2$ and $t^{\prime}=-0.15$ (b), $U=1$ and $t^{\prime}=-0.1$ (c), $U=0.5$ and $t^{\prime}=-0.015$ (d), all in two dimensions. The shaded regions correspond to phase separation: red (darker grey in the black and white version) to PS1, whereby one of the component phases is the half-filled EI state, and light blue (lighter grey) to PS2. Solid and dashed lines correspond to solutions of Eqs. (1516) and (13),(17) [PS1 and PS2 conditions, respectively]; bold vertical solid lines at $n=1$ denote a stable single-phase EI. Dotted line are the boundaries of the region around $n=1$, where the EM solution exists and has the lowest energy among the uniform single-phase states. Outside the phase separation region, the dashed-dotted lines show boundaries between single-band phases SB1 and SB2 (with the Fermi level within the broad or the narrow band, respectively) and semimetallic state (hatched).
@@
[…]
@@
Bold solid (dashed) lines at the borders of the PS1 (PS2) regions correspond to solutions to Eqs. (1516) [for PS2, Eqs. (13) and (17)]. Lighter lines correspond to continuations of these solutions within the phase separation regions; while these do not correspond to any transitions (see above, Sec. IV), they help visualise the overall structure of the phase diagram. We note that all four diagrams, which sweep different parameter regimes from strong to weak interaction, are remarkably similar in this regard. Therefore there is no doubt that results for other parameter values, or indeed for the three-dimensional case, would be qualitatively similar.
The parameter space outside the PS1 and PS2 regions is occupied, for the most part, by the single-band phases SB1 (with partially-filled broad band) and SB2 (whereby the chemical potential lies within the narrow band). In Fig. 7, the boundaries of the corresponding regions are shown with dashed-dotted lines. While the boundary at half-filling is continuous[25] (with increasing density, the broad band is filled at $n=1$ and subsequently the filling of the narrow band begins), no direct continuous transition between SB1 (with filled narrow band) and SB2 phases is possible in the $12>n>$ 1 region. In the large-$U$ case (see Fig. 7 a), the intervening area is entirely taken over by phase separation, PS2. In a completely filled system at $n=2$, the difference between SB1 and SB2 phases disappears, hence the boundaries of these two regions must meet (which is indeed the case for all parameter values used in Fig. 7). This occurs when the upper edges of the two filled Hartree bands coincide,
@@
Within our model, Eq. (1), the excitonic metal (EM) phase with $n\neq 1$ has a negative compressibility (see Sec. III) and does not appear on the phase diagram. Dotted lines within the phase separation regions in Fig. 7 show the areas around half-filling where the EM phase has the lowest energy among the uniform single-phase solutions. For weaker $U$ of Fig. 7 c,d the latter is always the case whenever the EM solution exists. For $U=2$, (see Fig. 7 c), there is a narrow region immediately above the upper (concave) boundary of the EM area at $n>1$, where the EM solution is present, yet has a higher energy than a single-band one. Within and around this region we sometimes encounter a situation where several EM solutions are present; the lowest-energy one has a negative compressibility. Finally, at $U=4$ this region extends upwards and to the right far into the larger-$E_{d}$,$n$ range, although in some cases the EM solution there may be spurious (corresponding to a local energy maximum). Elsewhere, the boundary of the EM region corresponds to a continuous transition ($\Delta\rightarrow 0$) into the lowest-energy uniform non-excitonic phase. We also note that the EM area significantly expands whenever the absolute value of $t^{\prime}$ is decreased. We will continue our discussion of the EM phase in the following section.
We recall that our selection of single phase states and component phases for phase separation includes uniform mean-field solutions only. We expect that this affects the validity of our results primarily at commensurate fillings, especially at $n=1$ beyond the region where the uniform EI phase is stable (which includes the charge/orbital ordering at small $|E_{d}|$, see Refs. 5, 6, 20). Possible instability of a uniform semimetal at a fractional filling with respect to charge/orbital modulation is another issue which falls beyond the scope of this work. Finally, when the narrow-band hopping $t^{\prime}$ is decreased below the critical value required to stabilise the uniform EI, the EI state at $n=1$ acquires spatial modulation of both charge density and excitonic correlations[6]. We expect that this should not significantly affect the behaviour in the doped regime at a small but finite $t^{\prime}$: within the overall picture described above, the modulated EI would take place of the uniform EI as a phase component in the PS1 region; apart from a quantitative change, the overall structure of the phase diagram presumably remains unaffected.

### v2

# Phys. Rev. B, in press. arXiv:2507.13967 Excitonic Insulator and the Extended Falicov–Kimball Model Away from Half-Filling

@@ Abstract
We consider an extended spinless Falicov–Kimball model at an arbitrary doping level, focusing on the range of parameter values where a uniform excitonic insulator is stabilised at half-filling. In the two-dimensional case, we compare the properties of possible uniform phases and construct the Hartree–Fock phase diagrams, which include sizeable phase separation regions. It is seen that the excitonic insulator can appear as a component phase in a mixed-phase state in a broad interval of doping levels. In addition, in a certain range of parameter values the excitonic metal (doped excitonic insulator) is identified as the lowest-energy uniform phase. We suggest that this phase, which is unstable with respect to phase separation, may be stabilised when the phase separation is suppressed by the long-range Coulomb interaction. Overall, we find that excitonic correlations can affect the behaviour of the system relatively far away from half-filling.
@@ II THE MODEL AND MEAN-FIELD EQUATIONS
The self-consistent mean-field equations for the excitonic phase are derived in a standard way (e.g., by diagonalising the decoupled Hamiltonian) and take form
@@
$n_{c}-n_{d}$ $=$ $\frac{1}{N}\sum_{\vec{k}}\frac{\left(\xi_{{\vec{k}}}+t^{\prime}\epsilon_{\vec{k}}\right)\left(n^{1}_{\vec{k}}-n^{2}_{\vec{k}}\right)}{\sqrt{(\xi_{{\vec{k}}}+t^{\prime}\epsilon_{\vec{k}})^{2}+4|U\Delta|^{2}}}\,$, (4)
which is valid at all values of bandfilling $0<n<2$,
@@
$\epsilon^{1,2}_{\vec{k}}$ $=$ $\frac{1}{2}\left[E_{d}+Un+\left(1+t^{\prime}\right)\epsilon_{k}\right]\mp$ (6)
@@
are quasiparticle energies in the two new bands, $\epsilon_{\vec{k}}=-\cos k_{x}-\cos k_{y}(-\cos k_{z})$ is the tight-binding dispersion in two (three) dimensions, and
$\xi_{\vec{k}}\equiv\xi(\epsilon_{\vec{k}})=E_{d}+U(n_{c}-n_{d})-\epsilon_{\vec{k}}\,.$ (7)
Generally, the net energy of the excitonic phase (“excitonic metal”, see Sec. III below) can be evaluated as [cf. Eq. (2)]
@@
At half-filling and at $T\rightarrow 0$, the quantity $n^{2}_{\vec{k}}$ vanishes, while $n^{1}_{\vec{k}}\equiv 1$ for all values of $\vec{k}$; this corresponds to an EI, and Eq. (8) then yields $E_{EI}$, the energy of EI. We shall assume [without loss of generality – see Eq. (3)] that the quantity $\Delta$ is real and positive.
@@
At smaller $U$, one also finds semimetal solutions with two partially-filled Hartree bands, centred around $Un_{d}$ (the broad band) and $E_{d}+Un_{c}=E_{d}+U(n-n_{d})$ (the narrow band). Introducing $\lambda_{c}=\mu-Un_{d}$, we then find for the energy difference $|t^{\prime}|\lambda_{d}$ between $\mu$ and the centre of the narrow band,
@@
$n_{d}=\int_{-d}^{\lambda_{d}}\nu(\epsilon)d\epsilon\,,\,\,\,n-n_{d}=\int_{-d}^{\lambda_{c}}\nu(\epsilon)d\epsilon\,$, (10)
@@
$E=E_{d}n_{d}+Un_{c}n_{d}+\int_{-d}^{\lambda_{c}}\epsilon\nu(\epsilon)d\epsilon+|t^{\prime}|\int_{-d}^{\lambda_{d}}\epsilon\nu(\epsilon)d\epsilon\,.$ (11)
The rest of the paper is concerned with numerical analysis of these solutions and of the ensuing phase diagrams in the interaction range $|t^{\prime}|\ll U/2d\stackrel{{\scriptstyle<}}{{\sim}}1$ (weak to moderately strong coupling). We shall be interested in the low-temperature limit, $T\rightarrow 0$ throughout. While we consider the two-dimensional (2D) case only (aside from the Appendix), it is expected that the mean-field analysis of a three-dimensional (3D) system would yield qualitatively similar results.
@@ III THE UNIFORM SOLUTIONS AND THEIR INSTABILITIES
We begin with the excitonic phase, which at half-filling, $n=1$, corresponds to the well-known excitonic insulator (EI) state, characterised by a non-zero value of the off-diagonal average $\Delta$ and a gap of the order of $U\Delta$ in the electron spectrum. Properties of EI phase and its stability have been extensively addressed in the literature[2, 3, 7]. Presently we find, that once the EI solution is present at $n=1$, a similar solution to the mean field equations, Eqs. (34), persists also in a certain range of doping values around $n=1$. The chemical potential then lies within the upper or lower hybridised band, away from the excitonic band gap, hence such a phase should be more properly called excitonic metal (EM). This term, which was used previously in the context of doped Mott insulators[17], is now more commonly applied to the case of doped and/or otherwise imperfect EI, with a non-zero density of states at the Fermi level[18, 20, 19].
Fig. 1 illustrates the typical behaviour of EM solution in the case of moderate (a) or weak (b) Coulomb repulsion $U$. Numerical calculations are somewhat tedious, in particular because even relatively small values of $t^{\prime}$ strongly affect the EM bandstructure and may, e.g., shift the minimum of the upper hybridised band away from $\vec{k}=0$. For better convergence, our computations here and in Figs. 4 and 7 below included a very small temperature smearing (corresponding to $T=10^{-5}$) in the equations (34) for the EM phase away from half-filling; this does not affect the results, which therefore represent the $T\rightarrow 0$ case.
@@
For $U=2$, the excitonic metal solution is found within the doping range $0.93<n<1.38$ and with increasing $n$ the value of $n_{d}$ increases from zero to one. Accordingly, at the two endpoints the excitonic solution merges with the two different single-band phases. At these endpoints, the off-diagonal average $\Delta$ shows square-root features; elsewhere, it has a maximum at $n=1$ and a smoother feature near $n=1.3$, the latter reflecting a feature in the bandstructure. In the weak-interaction regime of $U=0.5$, the excitonic metal solution arises for $0.91<n<1.12$, merging with the semimetal solution at the endpoints.
Importantly, everywhere away from half-filling the compressibility $\partial\mu/\partial n$ is negative (see Appendix), signalling an instability with respect to an inhomogeneity formation and ultimately to phase separation. The dependence $\mu(n)$ shows a jump at $n=1$. This point, where the value of $\mu$ lies within the gap between the lower (filled) and the upper (empty) bands [see Eq. (6)], has to be considered separately.
@@
We remark that analysis of a “pure” Falicov – Kimball model [Eq. (1) with $t^{\prime}=0$] yields qualitatively similar behaviour of mean-field solutions, although quantitative changes associated with the non-zero $t^{\prime}$ are in some cases appreciable. We include a finite $t^{\prime}<0$ in our analysis due to the peculiar features of the $t^{\prime}=0$ case. These include an instability of excitonic insulator at $n=1$ (for all values of $U$ and $E_{d}$), as clarified in Refs. [6, 7]. For the numerical data shown in the plots, we chose the values of $t^{\prime}$ well above the respective critical values required to stabilise the homogeneous excitonic insulator state at half-filling; these critical values correspond to a second-order transition into a spatially modulated phase[6, 7].
@@
where the values of $V_{0}$ and/or $V_{1}$ exceed certain numerically small critical values. While this leads to numerical changes, we tentatively verified that at the qualitative level, the conclusions of this section (that the excitonic metal state has a negative compressibility and also, typically, the lowest energy) remain valid. The important difference is that in the presence of a bare hybridisation one has to distinguish not between non-hybridised ($\Delta=0$) and hybridised (excitonic metal or insulator) solutions as above, but rather between phases with small and large $\Delta$. In the former case, $\Delta$ vanishes in the limit $V_{0,1}\rightarrow 0$, as the corresponding solution evolves into either single-band or semimetal one. This does not happen in the case of large-$\Delta$ solutions, corresponding to EM or EI (notwithstanding an instability at small $V_{0,1}$, which is similar to the instability at small $t^{\prime}$). Throughout the rest of the paper, we will be including the effects of $t^{\prime}$ only.
@@ IV PHASE SEPARATION
We note that the question of EI stability at half-filling has been addressed by a variety of methods (see, e.g., Refs. 5, 6, 15, 7), also beyond the simple mean-field approximation. Within the range of parameter values where the uniform EI is stable at $n=1$, this half-filled EI will therefore be stable also in a phase-separated state away from the (overall) half-filling, provided that the EI is in the thermodynamic equilibrium [see Eqs. (1617)] with the other constituent phase.
@@
At smaller $U$, this can preempt the EI – single-band PS1 phase separation and may also render the EI state at $n=1$ unstable (basically, the single-band state that would have been in thermodynamic equilibrium with the EI becomes unstable with respect to this second type of phase separation). We will now follow these two phase separation scenarios in a typical situation (the details may vary, depending on parameter values).
@@
With lowering $n$ further, the stability of the uniform phase $A$ is lost at a point $n=n_{A}^{*}$ where either $F_{A}$ or $F_{AB}$ vanishes (whichever occurs first). In the case where $F_{A}(n_{A}^{*})=0$ [while either $F_{AB}(n_{A}^{*})<0$ or Eq. (14) cannot be solved], lowering $n$ further leads to phase separation into phase A (with $n_{A}=n_{A}^{*}$) and the EI phase; the EI fraction expands until phase A disappears at half-filling ($n=1$) and the system turns into a uniform EI. At this point, the value of the chemical potential suffers a negative jump (within the EI spectral gap), and reducing $n$ further leads to a phase separation into EI and a single-band phase with empty narrow band.
@@
These two situations are exemplified by Fig. 6. For the values of parameters used in Fig. 6 (a), Eq. (16) is satisfied at $n<1.68$, and $F_{A}(n)$ changes sign at $n_{A}^{*}\approx 1.522$, signalling phase separation into a half-filled EI and the single-band phase A with $n=n_{A}^{*}>1$ (phase separation of the type PS1). On the other hand, the value of $F_{AB}$, which would describe phase separation into two different single-band phases, changes sign at $n\approx 1.492<n_{A}*$, and this latter scenario is therefore irrelevant (as the uniform phase A at this doping level is already unstable with respect to the other type of phase separation). In the case of Fig. 6 (b), Eq. (14) for those parameters values can be solved only in a very narrow doping range $1.608<n<1.612$ [cf. Fig. 3 (b), where the overall negative slope of the dashed line in the centre of the figure is very small]. The value of $F_{AB}(n)$ changes sign at $n_{A}^{*}\approx 1.611$, whereas $F_{A}(n)$ vanishes at $n\approx 1.610<n_{A}^{*}$ and the EI state is therefore irrelevant. Phase separation is thus of the PS2 type, corresponding to a mixture of two single band phases with $n=n_{A}^{*}$ and $n=n_{B}^{*}\approx 0.756$. We note that the precise boundary between the two scenarios depends also on the value of $t^{\prime}$, as increasing the latter tends to tilt the balance in favour of PS1.
@@ V THE PHASE DIAGRAMS
Figure 7: (colour online) Calculated phase diagrams for the 2D EFKM with $U=4$ and $t^{\prime}=-0.3$ (a), $U=2$ and $t^{\prime}=-0.15$ (b), $U=1$ and $t^{\prime}=-0.1$ (c), $U=0.5$ and $t^{\prime}=-0.015$ (d). The shaded regions correspond to phase separation: red (darker grey in the black and white version) to PS1, whereby one of the component phases is the half-filled EI state, and light blue (lighter grey) to PS2. Solid and dashed lines correspond to solutions of Eqs. (1617) and (14),(18) [PS1 and PS2 conditions, respectively]; bold vertical solid lines at $n=1$ denote a stable single-phase EI. Dotted lines are the boundaries of the region around $n=1$, where the EM solution exists and has the lowest energy among the uniform single-phase states. Outside the phase separation region, the dashed-dotted lines show boundaries between single-band phases SB1 and SB2 (with the Fermi level within the broad or the narrow band, respectively) and semimetallic state (hatched).
@@
While in the weakly-interacting case of Fig. 7 d this equality is satisfied (within our numerical accuracy) at the upper point of the EI line, $E_{d}\approx 1.525$, in case of the strong interaction (Fig. 7 a) this takes place at $E_{d}\approx 0.24$, near the middle of the EI line which extends up to $E_{d}\approx 0.43$. In the intermediate cases of Fig. 7 b and c, Eq. (20) is satisfied, respectively, at $E_{d}\approx 0.56$ and $E_{d}\approx 1.205$, while the upper end of EI line corresponds to $E_{d}\approx 0.695$ and $E_{d}\approx 1.215$.
@@
Bold solid (dashed) lines at the borders of the PS1 (PS2) regions correspond to solutions to Eqs. (1617) [for PS2, Eqs. (14) and (18)]. Lighter lines correspond to continuations of these solutions within the phase separation regions; while these do not correspond to any transitions (see above, Sec. IV), they help visualise the overall structure of the phase diagram. We note that all four diagrams, which sweep different parameter regimes from strong to weak interaction, are remarkably similar in this regard. Therefore there is no doubt that results for other parameter values would be qualitatively similar.
The parameter space outside the PS1 and PS2 regions is occupied, for the most part, by the single-band phases SB1 (with partially-filled broad band) and SB2 (whereby the chemical potential lies within the narrow band). In Fig. 7, the boundaries of the corresponding regions are shown with dashed-dotted lines. While the boundary at half-filling is continuous[25] (with increasing density, the broad band is filled at $n=1$ and subsequently the filling of the narrow band begins), no direct continuous transition between SB1 (with filled narrow band) and SB2 phases is possible in the $2>n>1$ region. In the large-$U$ case (see Fig. 7 a), the intervening area is entirely taken over by phase separation, PS2. In a completely filled system at $n=2$, the difference between SB1 and SB2 phases disappears, hence the boundaries of these two regions must meet (which is indeed the case for all parameter values used in Fig. 7). This occurs when the upper edges of the two filled Hartree bands coincide,
@@
Within our model, Eq. (1), the excitonic metal (EM) phase with $n\neq 1$ has a negative compressibility (see Sec. III and Appendix) and does not appear on the phase diagram. Dotted lines within the phase separation regions in Fig. 7 show the areas around half-filling where the EM phase has the lowest energy among the uniform single-phase solutions. For weaker $U$ of Fig. 7 c,d the latter is always the case whenever the EM solution exists. For $U=2$, (see Fig. 7 c), there is a narrow region immediately above the upper (concave) boundary of the EM area at $n>1$, where the EM solution is present, yet has a higher energy than a single-band one. Within and around this region we sometimes encounter a situation where several EM solutions are present; the lowest-energy one has a negative compressibility. Finally, at $U=4$ this region extends upwards and to the right far into the larger-$E_{d}$,$n$ range, although in some cases the EM solution there may be spurious (corresponding to a local energy maximum). Elsewhere, the boundary of the EM region corresponds to a continuous transition ($\Delta\rightarrow 0$) into the lowest-energy uniform non-excitonic phase. We also note that the EM area significantly expands whenever the absolute value of $t^{\prime}$ is decreased. We will continue our discussion of the EM phase in the following section.
We recall that our selection of single phase states and component phases for phase separation includes uniform mean-field solutions only. We expect that this affects the validity of our results primarily at commensurate fillings, especially at $n=1$ beyond the region where the uniform EI phase is stable (which includes the charge/orbital ordering at small $|E_{d}|$, see Refs. 5, 6, 15). Possible instability of a uniform semimetal at a fractional filling with respect to charge/orbital modulation is another issue which falls beyond the scope of this work. Finally, when the narrow-band hopping $t^{\prime}$ is decreased below the critical value required to stabilise the uniform EI, the EI state at $n=1$ acquires spatial modulation of both charge density and excitonic correlations[6]. We expect that this should not significantly affect the behaviour in the doped regime at a small but finite $t^{\prime}$: within the overall picture described above, the modulated EI would take place of the uniform EI as a phase component in the PS1 region; apart from a quantitative change, the overall structure of the phase diagram presumably remains unaffected.